OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. y + 2 = 1/3(x – 5); (–4, 3)

OpenStudy (anonymous):

y = –3x + 15 y = –3x – 9 y = –4x + 3 y = –1/3x – 11

OpenStudy (anonymous):

y=3x-9

OpenStudy (anonymous):

Can you help me with one more @Fireflame13 ? whether the lines for the pair of equations are parallel, perpendicular, or neither. y = –2/3x + 1 2x – 3y = –3

OpenStudy (anonymous):

*Tell whether

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

thank you so much! (:

OpenStudy (anonymous):

perpendicular, 1, and -3. 2xand single y.